# GaussianDWM: 3D Gaussian Driving World Model for Unified Scene Understanding and Multi-Modal Generation

### v1

@@ Appendix C Implementation Details
Our training pipeline consists of three stages. In the first stage, we train the Gaussian tokenizer, projector, and the proposed sampling strategy independently. We then integrate these components with the LLM and perform joint fine-tuning. For the LLM, we adopt a LoRA-based fine-tuning strategy. All models in this stage are trained using 16 NVIDIA A100 GPUs. In the second stage, we train the multi-modal generation module. We start by training a low-resolution ($224\times 400$) RGB video generation model, extend it to low-resolution RGB-D generation, and finally refine it into a high-resolution ($424\times 800$) RGB-D video generator using a mixed-frame-length strategy. The model is optimized using simulation-free rectified flow and a v-prediction loss [vprediction]. In the final stage, we perform end-to-end joint optimization over all components to ensure consistency between scene understanding and scene generation.
@@ Appendix D Ablation Study
GPT-4o [gpt4] - - - 0.223 0.244
LLaVA-OV [LLaVA] arXiv 24 7B - 0.221 0.284
OmniDrive [omnidrive] CVPR 25 7B 0.380 0.326 0.686
OmniDrive-2D [omnidrive] CVPR 25 7B 0.383 0.325 0.671
OmniDrive-BEV [omnidrive] CVPR 25 7B 0.356 0.278 0.595
@@
Table 6: OmniDrive dataset [omnidrive] Scene Understanding Results. We further evaluate our model on the OmniDrive dataset, comparing against strong baselines including GPT-4o [gpt4], LLaVA [LLaVA], OmniDrive [omnidrive], and Hermes [hermes]. Our GaussianDWM consistently achieves state-of-the-art performance across all evaluation metrics. These results again confirm the effectiveness of our 3D Gaussian scene representation and task-aware sampling strategy in enhancing LLM-based scene understanding under complex driving scenarios.
@@ Appendix E Extensive Experiments
OmniDrive dataset [omnidrive] Scene Understanding Results. We further evaluate our model on the OmniDrive dataset, comparing against strong baselines including GPT-4o [gpt4], LLaVA [LLaVA], OmniDrive [omnidrive], and Hermes [hermes]. Our GaussianDWM consistently achieves state-of-the-art performance across all evaluation metrics. These results again confirm the effectiveness of our 3D Gaussian scene representation and task-aware sampling strategy in enhancing LLM-based scene understanding under complex driving scenarios.

### v2

@@ Appendix C Implementation Details
Our training pipeline consists of three stages. In the first stage, we train the Gaussian tokenizer, projector, and the proposed sampling strategy independently. We then integrate these components with the LLM and perform joint fine-tuning. For the LLM, we adopt a LoRA-based fine-tuning strategy. All models in this stage are trained using 16 NVIDIA A100 GPUs. In the second stage, we train the multi-modal generation module. We start by training a low-resolution ($224\times 400$) RGB video generation model, extend it to low-resolution RGB-D generation, and finally refine it into a high-resolution ($424\times 800$) RGB-D video generator using a mixed-frame-length strategy. The model is optimized using simulation-free rectified flow and a v-prediction loss [15]. In the final stage, we perform end-to-end joint optimization over all components to ensure consistency between scene understanding and scene generation.
@@ Appendix D Ablation Study
GPT-4o [1] - - - 0.223 0.244
LLaVA-OV [32] arXiv 24 7B - 0.221 0.284
OmniDrive [46] CVPR 25 7B 0.380 0.326 0.686
OmniDrive-2D [46] CVPR 25 7B 0.383 0.325 0.671
OmniDrive-BEV [46] CVPR 25 7B 0.356 0.278 0.595
@@
Table 6: OmniDrive dataset [46] Scene Understanding Results. We further evaluate our model on the OmniDrive dataset, comparing against strong baselines including GPT-4o [1], LLaVA [32], OmniDrive [46], and Hermes [64]. Our GaussianDWM consistently achieves state-of-the-art performance across all evaluation metrics. These results again confirm the effectiveness of our 3D Gaussian scene representation and task-aware sampling strategy in enhancing LLM-based scene understanding under complex driving scenarios.
@@ Appendix E Extensive Experiments
OmniDrive dataset [46] Scene Understanding Results. We further evaluate our model on the OmniDrive dataset, comparing against strong baselines including GPT-4o [1], LLaVA [32], OmniDrive [46], and Hermes [64]. Our GaussianDWM consistently achieves state-of-the-art performance across all evaluation metrics. These results again confirm the effectiveness of our 3D Gaussian scene representation and task-aware sampling strategy in enhancing LLM-based scene understanding under complex driving scenarios.